The capacities of two trucks are 3 tons and 4 tons. If the larger truck makes 3 more trips than the smaller, its total haul is 20 more tons than the smaller one. How many trips does each truck make?
Found 2 solutions by Boreal, addingup:
Answer by Boreal(15235)

large truck takes x trips, smaller x-3 trips
4x-20=3(x-3)
4x-20=3x-9
x=11 trips larger truck and 44 tons.
x-3=8 trips smaller truck and 24 tons, 20 fewer.

Answer by addingup(3677)

4( x +3) = 3x + 20
4x + 12 = 3x + 20
x = 8 the small truck makes 8 trips and the large truck 8+3=11
--------------------------------------
Check:
4 tons * (8 + 3) = 4 * 11 = 44
(3 tons * 8) + 20 = 24 + 20 = 44
The answer is correct.